OpenStudy (albert0898):

How do you identify the independent and dependent variable? What is the independent and dependent variable? Give examples please!

OpenStudy (anonymous):

independent variable is the input, dependent is the output

OpenStudy (albert0898):

So how would you identify that on a coordinate grid?

OpenStudy (anonymous):

independent variable would be the length is ur _____, dependent variable is the pleasure of her org___

OpenStudy (anonymous):

independent: x (horizontal value) dependent: y (vertical value)

OpenStudy (anonymous):

In general, what is commonly used is expressions like y = f(x) = (some expression in x). "x" is the independent variable as you can select an x (you start with an x, that is) and you have to go through computation to arrive at a value of y. y depends on the selected x.

OpenStudy (anonymous):

If you started with th "y", the expression in x would not immediately produce the x. You would then have to redefine the equation in terms of x = f(y) = (some expression in y) and then y would be the independent variable. If you use the generally-accepted Cartesian coordinate system with y vertical, this last expression is a relation not an equation.

OpenStudy (albert0898):

Here's a picture of the problem...

OpenStudy (anonymous):

You have it identified correctly.

OpenStudy (albert0898):

Did I do everything correctly? And what does it mean to identify the independent and dependent variable? What should I write or do to identify it?

OpenStudy (anonymous):

There is some function where y = "tickets sold" = f(x) where x = the particular day. You don't have to create a function since the problem is not asking you to do so. Just be aware that some function can be made to fit the graph.

OpenStudy (anonymous):

So, simply, you could write "tickets sold" = f(day) and day is the independent variable.

OpenStudy (albert0898):

I don't quite understand your last statement

OpenStudy (anonymous):

Well it's a little abstract because we don't have an actual explicit function like x^2 or (day)^2 or 3 x day. But one does exist. One only has to realize that when one looks at the graph, he will look to the day first and then to the tickets sold for that day.

OpenStudy (albert0898):

I understand that, so what do I have to write for where it says 'identify the independent and dependent variable.'

OpenStudy (anonymous):

It's the "for that day" that gives the graph it's function character. As for indentifying the independent variable, that is "day" which is associated with "x" or "horizontal". The dependent variable is "tickets sold" which is associated with y or vertical.

OpenStudy (anonymous):

You might be trying to make the problem harder than it is. We all do that from time to time.

OpenStudy (albert0898):

Thanks for the help! :)

OpenStudy (anonymous):

you're quite welcome.